OpenStudy (anonymous):

3/0.8=12/b solve for b

OpenStudy (anonymous):

Cross multiply and bring b to the other side.

OpenStudy (anonymous):

I can guide you step by step if you want.

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

When you have two fractions like that, you need to take the denominators to the other side and cross multiply into the numerators. a/b = c/d will become a*d = c*d. What would it be for the above problem?